OpenStudy (anonymous):

Write the equation of the line perpendicular to 9x + 2y = 18 and passes through the point (1, 4).

OpenStudy (anonymous):

can you first find the slope the line given by: 9x+2y=18

OpenStudy (anonymous):

just wait ! i will get a pencil

OpenStudy (anonymous):

ohkay :) thanks

OpenStudy (anonymous):

9x + 2y = 18 => 2y = -9x + 18 => y = (-9/2)x + 9. Slope of this line = (-9/2)

OpenStudy (anonymous):

so we got the slope ! let us move to next part of the problem !

OpenStudy (anonymous):

Slope of the line ┴ to this line is -(1/(-9/2)) = 2/9

OpenStudy (anonymous):

now It passes thro' (1,4)

OpenStudy (anonymous):

Just replace x and y for 1 and 4: \[9 * 1 + 2 * 4 = 17\] The equation is : \[9x + 2y = 17\] The slope is given by the ratio of the two coefficients. The constant value can be found as shown above.

OpenStudy (anonymous):

so the equation is: => (y - 4) = (2/9) * (x - 1) => 9(y - 4) = 2(x - 1) => 9y - 36 = 2x - 2 => 2x - 9y = -34